src2tex version 2.12

#                                                                                       #

src2tex* version 2.12 (Sep 9, 1996)

Kazuo Amano† and Shinichi Nomoto‡

#                                                                                       #

#  Makefile  #

#  preliminaries  #

Here we just give preliminary definitions. If you don’t have gcc, please replace it with cc and plug suitable values into CFLAGS.

#                                                                                       #
CC      = gcc
CFLAGS  = -O2
SRCS    = fileio.c getdata.c langflag.c modflag.c options.c pas bold.c \
        tools.c text2tex.c src2tex.c
OBJS    = fileio.o getdata.o langflag.o modflag.o options.o pas bold.o \
        tools.o text2tex.o src2tex.o

#  make all  #

“make all” gives you src2tex and src2latex. Hopefully, you do not encounter any errors.

#                                                                                       #
all :
        $(CC) $(CFLAGS) -DUNIX -DTEX -c $(SRCS)
        $(CC) $(OBJS) -o src2tex
        $(CC) $(CFLAGS) -DUNIX -DLATEX -c $(SRCS)
        $(CC) $(OBJS) -o src2latex

#  make install  #

“make install” just strips src2tex and src2latex, and removes object files. You have to move executables to your favorite directory by yourself. No online manual is available at the moment.

#                                                                                       #
install :
        strip ./src2tex
        strip ./src2latex
        rm -f ./$(OBJS)

* Permission to use, copy, and modify this software and its documentation is granted under no conditions. However, the authors would be very happy if users could inform any modifications to kamano@tansei.cc.utokyo.ac.jp. Since src2tex is a free software, there is no warranty of any kind for the program.

† Department of Mathematics, Josai University, Keyakidai, Sakado, Saitama, 350-02, JAPAN (kamano@ po.iijnet.or.jp)

‡ Department of Mathematics, Josai University, Keyakidai, Sakado, Saitama, 350-02, JAPAN (m93093@ crescent.josai.ac.jp)

@-echo ""
@-echo " So, you have finished installation successively."
@-echo " Please copy src2tex and src2latex to a suitable"
@-echo " directory you like."
@-echo ""
install.man :
@-echo ""
@-echo " Sorry, there exists no online manual."
@-echo ""

#  make dos  #

If you are a DOS user, please just type “make dos” instead of “make all” and “make install”. Then you will get src2tex.exe and src2ltex.exe . IF you don’t have gcc, please replace it with cc or something like that.

#                                                                                       #
dos :
@-echo ""
@-echo " Please don't be panic when  \"make dos\"  does not work."
@-echo " What you have to do is to be patient and edit Makefile"
@-echo " carefully.  I am sure it would turn out to be a piece"
@-echo " of cake."
@-echo ""
$(CC) $(CFLAGS) -DDOS -DTEX -c $(SRCS)
$(CC) $(OBJS) -o src2tex
$(CC) $(CFLAGS) -DDOS -DLATEX -c $(SRCS)
$(CC) $(OBJS) -o src2ltex
del *.o

#  make verbose  #

“make verbose” generates noisy src2tex and src2latex. They report everything what they are doing. The authors are using them to find bugs.

#                                                                                       #
verbose :
$(CC) -DDEBUGGING -DUNIX -DTEX -c $(SRCS)
$(CC) $(OBJS) -o src2tex
$(CC) -DDEBUGGING -DUNIX -DLATEX -c $(SRCS)
$(CC) $(OBJS) -o src2latex

#  make plain  #

“make plain” generates plain src2tex and src2latex. They do not output any preambles and postambles; they simply output TEXt without anything like

\documentstyle⋯, \begindocument, \enddocument, \bye,

They would be useful when you have to insert program lists into your manuscript written in TEX.

#                                                                                       #
plain :
$(CC) -DPLAIN -DUNIX -DTEX -c $(SRCS)

$(CC) $(OBJS) -o src2tex
$(CC) -DPLAIN -DUNIX -DLATEX -c $(SRCS)
$(CC) $(OBJS) -o src2latex
clean :
rm -f ./$(OBJS)
rm -f ./a.out
rm -f ./src2tex
rm -f ./src2latex ./src2ltex
rm -f ./*~
rm -f ./*.bak
rm -f ./Makefile.tex
rm -f ./*.h.tex
rm -f ./*.c.tex
rm -f ./*.scm.tex
rm -f ./*.red.tex
rm -f ./contrib/src2tex2dvi.tex
rm -f ./contrib/src2tex2ps.tex
rm -f ./contrib/src2latex2dvi.tex
rm -f ./contrib/src2latex2ps.tex
rm -f ./contrib/c2tex.tex
rm -f ./contrib/*~
rm -f ./contrib/*.bak

#  make docs  #

“make docs” simply generates several examples by using src2tex2ps, src2latex2ps, etc.

#                                                                                       #
docs :
src2tex2dvi Makefile
src2tex2ps *.h
src2tex2ps *.c
src2latex2ps *.scm
src2tex2ps popgen.red
src2tex2ps -fourcolumns -pagelength 35 sqrt mat.red
cd ./contrib ; src2tex2dvi src2tex2dvi
cd ./contrib ; src2tex2dvi src2tex2ps
cd ./contrib ; src2tex2dvi src2latex2dvi
cd ./contrib ; src2tex2dvi src2latex2ps
cd ./contrib ; src2tex2dvi c2tex

#  make report  #

“make report” send your local /etc/motd file to one of the authors. This would be quite helpful for further development of src2tex.

#                                                                                       #
report :
@-if [ -f /etc/motd ] ; \
then \
        echo "TITLE:src2tex version 2.12" > make report.log; \
        echo USER:$(USER) >> make report.log; \
        echo HOSTNAME:$(HOSTNAME) >> make report.log; \
        cat /etc/motd >> make report.log; \

echo -n "reporting to kamano@po.iijnet.or.jp ... " ; \
mail kamano@po.iijnet.or.jp \
        < make report.log ; \
echo "done" ; \
rm make report.log ; \
else \
echo "make report failed" ; \
echo "thank you very much anyway" ; \
fi